# Co-Hopfianity is not a Profinite Property

### 04 Mar 2026 v1

@@ 1. Introduction
While some properties, such as being abelian, being nilpotent, satisfying a nontrivial law [4], and being polycyclic [22], are profinite properties, it is well known that many geometric and asymptotic properties are not. Notable examples of properties that fail to be profinite include amenability [15], Kazhdan’s property (T) [1], being conjugacy separable [9], Serre’s property FA [8], various finiteness properties [16], being torsion-free [7], left-orderability [7], and bi-orderability [13].
In this paper, we focus on the co-Hopfian property. A group is co-Hopfian if it does not contain a proper subgroup isomorphic to itself. This property can be viewed as a dual notion to the Hopfian property, which requires that every surjective homomorphism to itself is an isomorphism. While finitely generated residually finite groups are Hopfian (by Mal’cev [17]), the co-Hopfian property is far more delicate.
@@
Acknowledgment. Both authors thank Martin Bridson for fruitful discussions and for his interest in this work. We also thank Jan Kim for useful conversations. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2025-00513595).
@@ 2.1. Hopfian and co-Hopfian groups
As mentioned in the Introduction, the co-Hopfian property is more delicate than the Hopfian property. It is known that every hyperbolic group is Hopfian [24, Corollary 7.9], but a free group is not co-Hopfian. Hence, there are many hyperbolic groups that are not co-Hopfian. We will use the following result due to Z. Sela [23].
Theorem 2.1 ([23, Theorem 4.4]).
Every torsion-free one-ended hyperbolic group is co-Hopfian.
@@ 2.2. Profinite completion
be a short exact sequence where $Q$ is finitely presented. Then the map $\widehat{N}\to\widehat{G}$ induced by inclusion is an isomorphism if and only if $\widehat{Q}=1$ and $H_{2}(Q,\mathbb{Z})=0$.
@@ 2.4. A universal acyclic group
It is well established that hyperbolicity is not a profinite property. However, to the best of the authors’ knowledge, the question of whether toral relative hyperbolicity constitutes a profinite property has not been explicitly addressed in the existing literature. We note that, in the specific context of virtually compact special groups, toral relative hyperbolicity is indeed a profinite property, as detected by the profinite completion due to a recent result of P. Zalesskii [26, Theorem 3.7]. Our result demonstrates that this rigidity does not hold for general finitely generated residually finite groups.
@@ 3.1. The construction of a co-Hopfian group $G$
By Sela’s result, Theorem 2.1, a torsion-free one-ended hyperbolic group is co-Hopfian. The group $G$ is torsion-free hyperbolic by Theorem 2.7; it remains to show that $G$ is one-ended.

### 17 Mar 2026 v2

@@ 1. Introduction
While some properties, such as being abelian, being nilpotent, satisfying a nontrivial law [4], and being polycyclic [24], are profinite properties, it is well known that many geometric and asymptotic properties are not. Notable examples of properties that fail to be profinite include amenability [16], Kazhdan’s property (T) [1], being conjugacy separable [9], various finiteness properties [17], Serre’s property FA [8], bounded cohomology [12], being torsion-free [7], left-orderability [7], and bi-orderability [14].
In this paper, we focus on the co-Hopfian property. A group is co-Hopfian if it does not contain a proper subgroup isomorphic to itself. This property can be viewed as a dual notion to the Hopfian property, which requires that every surjective homomorphism to itself is an isomorphism. As finitely generated residually finite groups are Hopfian (by Mal’cev [18]), Hopficity is vacuously a profinite invariant among them. However, if we take the definition of a profinite invariant to just mean “detectable by finite quotients”, so that it does not a priori assume residual finiteness, we can find easy counterexamples: one can take any finitely generated non-Hopfian group and its largest residually finite quotient, like the Baumslag–Solitar group $\mathrm{BS}(2,3)$ and $\mathbb{Z}[1/6]\rtimes_{2/3}\mathbb{Z}$. While perhaps not that interesting, it is still worth mentioning. In contrast, the co-Hopfian property is far more delicate.
@@
Acknowledgment. Both authors thank Martin Bridson for fruitful discussions and for his interest in this work. We also thank Jan Kim and Francesco Fournier-Facio for useful discussions and suggestions which improved the exposition of the paper. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2025-00513595).
@@ 2.1. Hopfian and co-Hopfian groups
As mentioned in the Introduction, the co-Hopfian property is more delicate than the Hopfian property. It is known that every hyperbolic group is Hopfian (the torsion-free case was previously obtained by Sela [26], and the general case by Weidmann–Reinfeldt [27, Corollary 7.9]), but a free group is not co-Hopfian. Hence, there are many hyperbolic groups that are not co-Hopfian. We will use the following result, which was originally proved by Z. Sela for the torsion-free case [25, Theorem 4.4] and extended to all one-ended hyperbolic groups without the torsion-free assumption by C. Moioli [19].
Theorem 2.1 ([25, Theorem 4.4], [19]).
Every one-ended hyperbolic group is co-Hopfian.
@@ 2.2. Profinite completion
be a short exact sequence where $Q$ is finitely presented. Then the map $\widehat{N}\to\widehat{G}$ induced by inclusion is an isomorphism if $\widehat{Q}=1$ and $H_{2}(Q,\mathbb{Z})=0$.
@@ 2.4. A universal acyclic group
It is well established that hyperbolicity is not a profinite property. However, to the best of the authors’ knowledge, the question of whether toral relative hyperbolicity constitutes a profinite property has not been explicitly addressed in the existing literature. We note that, in the specific context of virtually compact special groups, toral relative hyperbolicity is indeed a profinite property, as detected by the profinite completion due to a recent result of P. Zalesskii [29, Theorem 3.7]. Our result demonstrates that this rigidity does not hold for general finitely generated residually finite groups.
@@ 3.1. The construction of a co-Hopfian group $G$
By Theorem 2.1, a one-ended hyperbolic group is co-Hopfian. The group $G$ is a torsion-free hyperbolic group by Theorem 2.7; it remains to show that $G$ is one-ended.
@@ 3.4. Profinite completions and finiteness properties
In our construction, $G$ is hyperbolic and hence finitely presented. However, we do not know whether $H$ is finitely presented. Although we obtain a finitely generated example, it remains open whether such an example can be finitely presented. We therefore pose the following question.
Question 3.8.
Is co-Hopfianity a profinite property among finitely presented residually finite groups?